OpenStudy (anonymous):

(2.55x10^-2) divided by (1.5x10^-14)

OpenStudy (anonymous):

Use scientific notation

OpenStudy (lalaly):

\[\large{\frac{2.55}{1.5} \times 10^{-2-14}}\]

OpenStudy (lalaly):

oopps sorry its 10 ^ { -2-(-14)} = 10^12

OpenStudy (lalaly):

\[\huge{1.7 \times 10^{12}}\]